# EmoDynamiX: Emotional Support Dialogue Strategy Prediction by Modelling MiXed Emotions and Discourse Dynamics

### v1

@@ 5.1 Experimental Setups
Datasets To make our model learn strategies that are beneficial for both social and task-oriented goals, we select two ESC datasets (in English) where dialogues have been human-evaluated and filtered to ensure that the conversational outcomes are positive and the strategies applied are socially appropriate: (i) ESConv (Liu et al., 2021), an emotional support dataset annotated by trained crowd-workers. It comprises 1,300 long dialogues, averaging 29.5 speaker turns per dialogue, and features 8 dialogue strategies. To ensure fair comparisons, we adhere to the official train/dev/test split [^1] https://huggingface.co/datasets/thu-coai/esconv; (ii) AnnoMI (Wu et al., 2022), an expert-annotated counselling dataset. It includes 133 long conversations, averaging 72.9 speaker turns per dialogue, and features 9 therapist strategies. We split this dataset randomly into train/dev/test sets in an 8:1:1 ratio. Detailed dataset statistics are provided in Appendix A.
Baselines To provide extensive comparisons, we choose baselines from three criteria: (i) SOTA LLMs for dialogue tasks with few-shot prompting: ChatGPT (OpenAI, 2023) and LLaMA3-70B (Meta, 2024); (ii) models fine-tuned on both next strategy prediction and utterance generation: BlenderBot-Joint (Liu et al., 2021); (iii) models fine-tuned for emotional support dialogue strategy prediction: MISC (Tu et al., 2022) and KEMI (Deng et al., 2023). Detailed descriptions can be found in Appendix E.
Implementation Details We implemented our proposed method using PyTorch (Paszke et al., 2019), initializing with the pre-trained weights from RoBERTa and employing the tokenization tools from Huggingface Transformers (Wolf et al., 2020). For optimization, we used the AdamW optimizer (Loshchilov and Hutter, 2019). Detailed hyperparameter settings can be found in Appendix F.
Model ESConv AnnoMI
[…]
LLaMA3-70B 15.36 18.45 8.38 9.01
+ 2 shot (random) 14.68 17.74 9.45 9.59
+ 2 shot (extractive) 17.70 21.47 9.52 9.13
ChatGPT 18.14 20.27 20.31 18.12
+ 2 shot (random) 17.24 19.91 17.02 15.72
+ 2 shot (extractive) 16.55 20.01 15.29 14.22
MISC 19.89 - - -
BlenderBot-Joint 20.91 24.50 17.90 25.10
KEMI 24.66 26.80 - -
EmoDynamiX ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$
Table 1: Experimental results on two ESC datasets. $\mathcal{M}$ stands for Macro F1 and $\mathcal{W}$ for Weighted F1 (same for Table 2). $\spadesuit$ indicates results reported in Deng et al. (2023) and all other results are produced by ourselves. † indicates statistically significant improvement ($p<0.05$) over the best baseline.
@@ 5.2 Overall Performance
As shown in Table 1, EmoDynamiX significantly outperforms all baseline models across all metrics on both datasets. Compared to the state-of-the-art joint strategy learning model KEMI on ESConv, our method delivers more diverse (Macro F1) and accurate (Weighted F1) predictions. While KEMI leverages domain-specific actual case knowledge from HEAL Welivita and Pu (2022) to better model user emotion states compared to MISC, which relies solely on commonsense knowledge from COMET, it still falls short in capturing fine-grained emotion transitions in dialogues. Our ERC-based mixed-emotion approach addresses this gap more effectively. Furthermore, by modelling the interactions between user emotion states and historical strategies using a heterogeneous graph, our model can predict strategies that are more consistent with the overall progress of the counseling process.
Comparisons with two LLM prompting baselines reveal that using LLMs alone for emotional support dialogue prediction is severely limited by their inherent biases, even with provided examples. More analysis on the impact of few-shot learning is provided in Appendix H Extensive experiments on AnnoMI further confirm that our method excels in more professional and low-resource settings. While LLM-based methods and joint learning methods exhibit a significant drop in performance (particularly LLaMA3-70B, which failed to discern these strategies and resorted to using only a few), our method maintains its advantage, showcasing its robustness and effectiveness.
Model ESConv AnnoMI
[…]
EmoDynamiX 27.40 30.81 27.48 33.09
Flattened Context 25.04↓2.36 28.68↓2.13 22.25↓5.23 27.25↓5.84
w/o Graph 25.72↓1.68 29.31↓1.50 26.95↓0.53 29.46↓3.63
w/o Mixed Emotion 25.90↓1.50 29.45↓1.36 24.71↓2.77 30.25↓2.84
w/o Discourse Parser 26.64↓0.76 30.12↓0.69 27.04↓0.44 31.59↓1.50
w/o Dummy Node 26.67↓0.73 30.16↓0.65 26.42↓1.06 29.21↓3.88
Table 2: Evaluation results of ablation study.
Model $\mathcal{R}$ $\mathcal{S}$ $\mathcal{A}$
EmoDynamiX 14.80 20.12 16.69
Flattened Context 14.04 18.90 10.65
w/o Graph 11.3 17.79 8.98
Table 3: Ablation study on three emotion-related strategies of ESConv.
@@ 5.3 Ablation Study
We conducted ablation studies on the two datasets, as shown in Table 2. Our analysis is organized around several research questions:
@@
RQ 1: Is Emotion & Strategy beneficial for strategy prediction? We compared two simplified versions based solely on the semantic modeling module with a classification head. The first version uses only flattened dialogue history with speaker roles inserted as tags (Flattened Context). This resulted in a significant drop in both Macro F1 and Weighted F1 scores compared to EmoDynamiX. We then augmented Flattened Context with user emotions and system strategies, also inserted as tags (w/o Graph). Although this augmentation led to some improvements over Flattened Context, there remains a significant performance gap compared to EmoDynamiX. This indicates that while incorporating emotions and strategies is beneficial, the effectiveness is still limited without our heterogeneous graph.
RQ 2: Does Mixed-emotion bring better results? To answer this question, we tested the effectiveness of our mixed emotion module by embedding user emotional states with one-hot vectors, omitting the mixed-emotion module (w/o Mixed Emotion). The resulting decreases in both metrics highlight the importance of leveraging emotion distributions, not just labels, for capturing fine-grained user emotion states. It is worth noting that mixed-emotion plays a more crucial role in low-resource settings, as evidenced by the more significant performance decline on AnnoMI.
@@
RQ 3: Do we need discourse structure? We excluded the discourse parser, connecting the nodes in simple sequential order (w/o Discourse Parser). Although this led to drops in both metrics, the declines were not substantial. We hypothesize that the domain gap between the discourse parsing dataset and the ESC datasets may limit the discourse parsing module’s effectiveness.
@@
RQ 4: Are dummy nodes better for information aggregation? We removed the dummy node and used traditional graph readout functions by concatenating mean and max node embeddings (Joshi et al., 2021). We observed performance decreases on both datasets, with a more significant decline on AnnoMI, indicating that our dummy node design is particularly beneficial in low-resource settings.
RQ 5: How does graph learning improve our model’s emotional intelligence? We compared three emotion-related dialogue strategies in ESConv: Reflection of Feelings ($\mathcal{R}$), Self-disclosure ($\mathcal{S}$), and Affirmation and Reassurance ($\mathcal{A}$). As observed in Table 3, incorporating user emotions and previous system strategies simply as tags into the flattened context (w/o Graph) is ineffective, as it drastically decreases performance across all three emotion-related strategies compared to Flattended Context. This finding highlights that modeling conversational dynamics with heterogeneous graphs can effectively mitigate preference bias and enhance our model’s emotional intelligence.
5.4 Case Study
@@
We illustrated a case study using a conversation snippet from the ESConv test set, as shown in Figure 3. The case involves the agent deciding which strategy to apply after the user’s emotional state has transitioned positively from Frustration (as Frustration is not a category in DailyDialog, our mixed-emotion module models it as moderated sadness with a little anger) to Joy due to the agent’s previous applied strategies Providing Suggestions and Self-disclosure. The ground truth here is to apply Affirmation and Reassurance to acknowledge the user’s positive transition and encourage the user to consolidate this positive mood. Our model made the correct prediction, and the visualization offers insights into the decision-making details. By looking at the attention weights of dummy node edges, we can observe the contribution of each node. In the early stages (shallow RGAT layer), the dummy node pays more attention to previously applied socio-emotional strategies. As the RGAT layer deepens, the dummy node starts to shift its attention to user’s emotional transition, with higher weights applied to edges connecting with the emotion state nodes (especially Frustration). In short, our graph-learning module effectively captures clues from emotion/strategy dynamics for better prediction. Additionally, the attention weights provide a certain extent of transparency to the previously "black-box" decision process.
6 Error Analysis
While our method achieved SOTA results, most of the predicted strategy labels still do not match the ground truths. We examined instances where our model failed to predict the desired strategies to provide directions for future improvement.
We discovered that correctly predicting emotion-related dialogue strategies is the most difficult challenge. For our performance on ESConv (confusion matrix in Appendix I), three emotion-related strategies (Reflection of feelings, Self-disclosure, and Affirmation and Reassurance) have an average F1 score of 22.14, which is significantly lower than the overall Macro F1 score of 27.40. Among the samples where emotion-related strategies were the true label but were wrongly predicted, 63% were categorized under content-providing strategies (Question, Providing Suggestions, Information, and Restatement or Paraphrasing) and Others. This indicates that the model struggles to choose the correct timing for pushing towards task-oriented goals, failing to recognize when users need emotional soothing rather than practical suggestions or information. If emotion-related strategies are not delivered when needed, the social-oriented goals of the dialogue are underachieved, harming the overall outcome of emotionally supportive dialogue. Therefore, it is crucial for future work to deepen the understanding of users’ fine-grained and dynamic emotional states to deliver strategies at the correct timing.
@@ 7 Conclusions
In this paper, we propose EmoDynamiX, a novel socio-emotional dialogue strategy predictor that uses heterogeneous graphs to model the conversational dynamics of user emotional states and previously applied system strategies. By using pre-trained expert models, we use the discourse structure and model the emotional states of the user in a fine-grained style. Additionally, we enable speaker role-aware information aggregation with a dummy node design. Our approach has demonstrated its effectiveness by significantly improving all baselines on two public ESC datasets. We hope that our dedicated model can be integrated into future LLM-based ESC agents to mitigate the inherent limitations of LLMs in strategy planning, ultimately providing better emotional support with better transparency.
@@
Limitations on Ground Truths Although the ESC datasets we use are all human-evaluated, we cannot fully ensure that no strategy other than the ground truth could have worked in the same context. However, we currently lack a reliable method to accurately model user preferences and exhaustively compare all possible strategies at each step. We believe that developing such effective user simulations is a promising direction for future work, which would enable us to identify optimal strategies for each step and even allow for multi-step planning beyond just predicting the next strategy.
More on Conversational Dynamics Our method focused solely on the socio-emotional strategies of the agent and the emotional states of the user. However, integrating the agent’s emotions and the user’s dialogue acts could enhance the system further. This compound state design might necessitate a new architecture, which we plan to explore in future work.
@@
Limitations on Expert Modules Since Emotion Recognition in Conversation (ERC) and discourse parsing are not the primary contributions of our research, we did not investigate the impact of using different model architectures or datasets. The training and integration of a cross-domain ERC module and discourse parser will be considered in future studies.
@@
Intent of Technology We insist that conversational AI should not be developed to replace humans. Therefore, it is crucial to maintain a clear line between AI and humans (Collectif et al., 2021). Given that the training data for conversational AI, including the two datasets we selected, are primarily curated by humans, the AI (especially those trained using end-to-end methods) may exhibit human-like behaviors. For instance, in Figure 1, the system learns to utilize Self-disclosure strategy by expressing feelings of loneliness, which is not consistent with ethical recommendations. We stress that strategies leading to such human-like behaviors should be applied cautiously and potentially restricted in real-world applications to ensure safety. We believe that our next strategy prediction approach, which allows us to explicitly set desired behaviors for AI, can provide better control over conversational AI in the future.
@@ Appendix A Datasets
STAC (Asher et al., 2016) is a multi-party dialogue corpus collected from an online game. It contains 1,081 dialogues, with an average of 8.5 speaker turns per dialogue. STAC includes 16 discourse dependency categories: Comment, Clarification Question, Elaboration, Acknowledgment, Continuation, Explanation, Conditional, Question-Answer Pair, Alternation, Question-Elaboration, Result, Background, Narration, Correction, Parallel, and Contrast.
DailyDialog (Li et al., 2017) is a Emotion Recognition in Conversations (ERC) dataset collected from an English learning website, its topics are closer to everyday issues and thus better suited for ESC compared to other popular counterparts collected from TV shows (Poria et al., 2019; Zahiri and Choi, 2018) or actor performances (Busso et al., 2008). DailyDialog includes 13,118 multi-turn dialogues, with an average of 7.9 speaker turns per dialogue. The emotion labels in this dataset encompass Ekman’s six basic emotions (Anger, Disgust, Fear, Joy, Sadness, Surprise) and a Neutral class.
Appendix B Implementation Details for Submodules
Appendix C Definitions of Strategies in ESConv
@@
Appendix D Definitions of Strategies in AnnoMI
@@
Appendix E Baselines
ChatGPT (OpenAI, 2023) is an advanced language model with 175 billion parameters developed by OpenAI. Using reinforcement learning from human feedback (RLHF), it generates human-like text and excels in natural language processing tasks such as conversation and content creation. The prompting is constructed with the template in Figure 5 and corresponding strategy definitions in Appendix C or Appendix D. To facilitate few-shot learning, we built a case library using the training data and extracted two examples from this library for each inference on the test data. We employed two methods for example extraction: random selection and selection based on similarity scores computed from sentence-BERT (Reimers and Gurevych, 2019) embeddings.
LLaMA3-70B (Meta, 2024) is an instruction-tuned language model with 70 billion parameters developed by Meta. Optimized for dialogue use cases, it outperforms many available open-source chat models on common industry benchmarks. The prompting templates and few-shot approach are the same as described above.
BlenderBot-Joint (Liu et al., 2021) is based on BlenderBot (Roller et al., 2021). It jointly predicts strategies and generates responses.

### v2

@@ 5.1 Experimental Setups
Datasets To make our model learn strategies beneficial for both social and task-oriented goals, we select two ESC datasets (in English) where dialogues have been human-evaluated and filtered to ensure that the conversational outcomes are positive and the strategies applied are socially appropriate: (i) ESConv (Liu et al., 2021), an ESC dataset annotated by trained crowd-workers. It comprises 1,300 dialogues and features 8 dialogue strategies. We adhere to the official train/dev/test split [^2] https://huggingface.co/datasets/thu-coai/esconv.; (ii) AnnoMI (Wu et al., 2022), an expert-annotated counselling dataset. It includes 133 dialogues and features 9 therapist strategies. We make the train/dev/test split with an 8:1:1 ratio. Detailed dataset statistics are provided in Appendix A.1. For both datasets, we set the context window size to 5 utterances, resulting in 18,376 samples for ESConv and 4,442 samples for AnnoMI.
Baselines To provide extensive comparisons, we choose baselines from three criteria: (i) prompting LLMs SOTA in dialogue tasks, using task description and supplementary information: ChatGPT (OpenAI, 2023) and LLaMA3-70B (Meta, 2024) with 2-shot learning (+2 shot) or emotion labels (+ ERC). We excluded Chain-of-Thought (Wei et al., 2024) prompting because it has already been shown to be ineffective for our task (Kang et al., 2024); (ii) fine-tuning LLMs as general-purpose dialogue strategy predictors: RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020) and LLaMA3-8B (Meta, 2024); (iii) specialized models for emotional support dialogue strategy prediction: MISC (Tu et al., 2022), MultiESC (Cheng et al., 2022), KEMI (Deng et al., 2023) and TransESC (Zhao et al., 2023). For more details see Appendix D.
Evaluation Metrics We use macro F1 score ($\mathcal{M}$-F1) and weighted F1 score ($\mathcal{W}$-F1) as metrics for evaluating the proficiency of strategy prediction models, since the ground truth strategies in the two datasets have been validated by human evaluators. Given the unbalanced nature of ESC datasets, the accuracy score is not an ideal choice as it can unfairly favor models that predominantly predict the majority classes. We also incorporate the preference bias score ($\mathcal{B}$) as defined by Kang et al. (2024) to quantify the extent to which the model favors its preferred strategies over non-preferred ones (implementation details in Appendix E). An ideal dialogue strategy predictor should achieve strong F1 scores while minimizing preference bias.
Implementation Details We implemented our proposed method using PyTorch (Paszke et al., 2019), initializing with the pre-trained weights from RoBERTa and employing the tokenization tools from Huggingface Transformers (Wolf et al., 2020). For optimization, we used the AdamW optimizer (Loshchilov and Hutter, 2019). Detailed hyperparameter settings can be found in Appendix H.
@@ 5.2 Overall Performance
As shown in Table 1, EmoDynamiX outperforms previous SOTA methods across all evaluation metrics. Compared to TransESC, EmoDynamiX significantly reduces the preference bias score by 38% and achieves higher F1 scores. This suggests that while TransESC also models dialogue state transitions, our ERC-based mixed-emotion approach captures the nuances of user emotion states more effectively, leading to better predictions, as further validated by our ablation study. Furthermore, compared to MultiESC, the previous SOTA model with a low bias score, EmoDynamiX excels across all metrics by a substantial margin. Another interesting result is that models based on specific case knowledge (KEMI) and general commonsense knowledge (MISC) are less effective as dialogue strategy predictors.
Comparisons with two LLM prompting baselines indicate that using LLMs alone for emotional support dialogue prediction is significantly constrained by their inherent biases, even when examples or emotion recognition is provided through prompts. The bias scores for LLM-prompting baselines are significantly higher, ranging from 0.77 to 1.39. Among the LLM-fine-tuning baselines, LLaMA3-8B achieves the highest F1 scores. RoBERTa generally exhibits a lower bias score, while BART emerges as a balanced option. Nevertheless, these baselines still lag behind EmoDynamiX by a considerable margin.
@@
Figure 4: Analysis on the correlation between the top-10 disagreement patterns (Ground Truth -> Prediction) and their most influential emotion categories.
@@ 5.3 Ablation Study
We conducted ablation studies (Table 2) and investigated whether the following modules improves the results of next strategy prediction:
@@
Modelling user emotions and agent strategies in dialogue context. We compared two simplified versions: (1) flattened dialogue history only (RoBERTa in Table 1); (2) flattened context with emotions and strategies inserted as tags (w/o Graph). Although w/o Graph outperforms RoBERTa, there remains a significant gap compared to EmoDynamiX. This indicates that while incorporating emotions and strategies is beneficial, the effectiveness is still limited without our graph-learning module.
Modelling mixed emotions. We modelled user emotional states with one-hot vectors instead (w/o Mixed Emotion). The resulting decreases in all metrics highlight the importance of leveraging emotion distributions, not just labels, for capturing fine-grained user emotion states.
@@
Modelling discourse structure. We connected the nodes in simple sequential order instead of discourse structure (w/o Discourse Parser). Although this led to drops in all metrics, the declines were not substantial. We hypothesize that the domain gap between the STAC dataset and the ESC datasets may limit the discourse parsing module’s effectiveness.
@@
Use of dummy nodes for information aggregation We replaced dummy nodes with traditional mean-max pooling (Joshi et al., 2021) (w/o Dummy Node). We observed performance decreases, with a more significant decline on AnnoMI, indicating that our dummy node design is particularly beneficial in low-resource settings.
6 In-depth Analysis of EmoDynamiX
@@
We illustrated a case study using a snippet from ESConv, as shown in Figure 3. The case involves the agent deciding which strategy to apply after the user’s emotional state has transitioned positively from Frustration (as Frustration is not a category in DailyDialog, our mixed-emotion module models it as moderated sadness with a little anger) to Joy. The ground truth here is Affirmation and Reassurance, which acknowledges the user’s positive transition and encourages consolidation of positive mood. By looking at the attention weights of dummy node edges, we can observe the contribution of each node in decision-making. We notice that, as the RGAT layer deepens, the dummy node shifts its attention from previously applied strategies to the user’s emotional transition, with higher weights applied to edges connecting with the emotion state nodes (especially Frustration). In short, our graph-learning module effectively captures clues from emotion/strategy dynamics.
We further study the disagreements between the predictions and human strategies using the confusion matrix (Appendix J). Determining the appropriate timing for emotion-related dialogue strategies versus task-oriented ones is particularly challenging. The model frequently predicts Providing Suggestions whereas the human strategies are among the three emotion-related ones: Reflection of feelings, Self-disclosure, and Affirmation and Reassurance. Notably, over 60% of emotion-related strategies were categorized as task-oriented ones, highlighting the difficulty in making choices between these two strategy categories, as also discussed by Galland et al. (2022). We further analyze the disagreement patterns between predicted and human strategies. By looking at their correlations with the primary emotion categories of emotion nodes with the highest attention weights (Figure 4), we find that "Neutral" contributes a larger proportion to these disagreements in general compared to its overall representation in the ERC module’s output distribution (59.22%, as shown in Appendix F). This suggests that strategies are easier to predict when the model can pick up the user emotions expressed in the context.
@@ 7 Conclusions
In this paper, we propose EmoDynamiX, a socio-emotional dialogue strategy prediction framework that aggregates expert models and uses heterogeneous graphs to model the conversational dynamics of user states and system strategies. Our approach significantly improves all baselines on two public ESC datasets, and takes a step towards transparency by analyzing attention weights in the in-depth study.
@@
Limitations on Ground Truths Although the ESC datasets we use have all been evaluated by humans, we cannot fully ensure that no other strategies, aside from the ground truth, could have been effective in the same context. However, we currently lack a protocol for human evaluation at the strategy-planning stage. Besides, human evaluation is more suitable to be performed after the generation stage.
@@
Limitations on Expert Modules Since ERC and discourse parsing are not the primary contributions of our research, we did not investigate the impact of using different model architectures or datasets for their training. The training and integration of a cross-domain ERC module and discourse parser could be considered in future studies.
@@
Intent of Technology We insist that conversational AI should not be developed to replace humans. Therefore, it is crucial to maintain a clear line between AI and humans (Ethique et al., 2024). Given that the training data for conversational AI, including the two datasets we selected, are primarily curated by humans, the AI (especially those trained using end-to-end methods) may exhibit human-like behaviors. For instance, in Figure 1, the system learns to utilize Self-disclosure strategy by expressing feelings of loneliness, which is not consistent with ethical recommendations. We stress that strategies leading to such human-like behaviors should be applied cautiously and potentially restricted in real-world applications to ensure safety. We believe that our next strategy prediction approach, which allows us to explicitly set desired behaviors for AI, can provide better control over conversational AI in the future.
@@ Appendix A Datasets
Figure 5: Strategy distributions of ESConv (left) and AnnoMI (right).
A.1 ESC Datasets
Following are the two public ESC datasets we use to evaluate our method.
@@ A.1 ESC Datasets
A.2 Datasets for Pre-training Expert Models
STAC (Asher et al., 2016) is used for pre-training our discourse parser. It is a multi-party dialogue corpus collected from an online game. It contains 1,081 dialogues, with an average of 8.5 speaker turns per dialogue. STAC includes 16 discourse dependency categories: Comment, Clarification Question, Elaboration, Acknowledgment, Continuation, Explanation, Conditional, Question-Answer Pair, Alternation, Question-Elaboration, Result, Background, Narration, Correction, Parallel, and Contrast.
DailyDialog (Li et al., 2017) is used to pre-train our emotion recognition module. It is an ERC dataset collected from an English learning website. Its topics are closer to everyday issues and thus better suited for ESC compared to other popular counterparts collected from TV shows (Poria et al., 2019; Zahiri and Choi, 2018) or actor performances (Busso et al., 2008). DailyDialog includes 13,118 multi-turn dialogues, with an average of 7.9 speaker turns per dialogue. The emotion labels in this dataset encompass Ekman’s six basic emotions (Anger, Disgust, Fear, Joy, Sadness, Surprise) and a Neutral class.
Appendix B Definitions of Strategies in ESConv
@@
Appendix C Definitions of Strategies in AnnoMI
@@
Appendix D Baselines
ChatGPT (OpenAI, 2023) is an advanced language model with 175 billion parameters developed by OpenAI. Using reinforcement learning from human feedback (RLHF), it generates human-like text and excels in natural language processing tasks such as conversation and content creation. The prompting is constructed with the template in Figure 6 and corresponding strategy definitions in Appendix B or Appendix C. To facilitate few-shot learning, we built a case library using the training data and extracted two examples from this library for each inference on the test data. We employed example extraction based on similarity scores computed from sentence-BERT (Reimers and Gurevych, 2019) embeddings.
LLaMA3 (70B & 8B) (Meta, 2024) is a series of instruction-tuned language models developed by Meta, featuring parameter sizes ranging from 8 to 70 billion. These models are specifically optimized for dialogue use cases and demonstrate superior performance compared to many existing open-source chat models on standard industry benchmarks. For the 70B variant, the prompting templates and few-shot methodologies align with those previously described. In contrast, the 8B variant uses flattened dialogue context with speaker tags as input (same for RoBERTa and BART) and employs LoRA (Hu et al., 2021a) for parameter-efficient fine-tuning.
RoBERTa (Liu et al., 2019) is a transformer-based language model that enhances BERT by using more training data, larger batch sizes, and dynamic masking, while removing the Next Sentence Prediction objective. Its robust training approach makes it more effective than the original BERT model across multiple benchmarks.
BART (Lewis et al., 2020) is a sequence-to-sequence model that combines a bidirectional encoder and an autoregressive decoder, effectively blending BERT and GPT architectures. It’s trained to reconstruct original text from corrupted input, making it highly versatile for tasks like text generation, summarization, and translation.